Suzy and Charolette are involved in a service project helping ederly neighbors fix up their yards for summer. If it takes Suzy 9 hours to do the job on her own, and it takes them 4 hours working together, how long will it take Charolette to fix up the yard by herself?
Answer by checkley75(3666)

9Y/(9+Y)=4 NOW CROSS MULTIPLY.
9Y=4(9+Y)
9Y=36+4Y
9Y-4Y=36
5Y=36
Y=36/5
Y=7.2 HOURS FOR CHARLOTTE TO DO THE JOB ALONE.
PROOF
9*7.2/(9+7.2)=4
64.8/16.2=4
4=4